# On Creating a Causally Grounded Usable Rating Method for Assessing the Robustness of Foundation Models Supporting Time Series

### v2

@@ 4 Solution Approach
Our solution approach consists of the following components: (1) Three perturbations that are applied to both numerical time series data and line plots. (2) Metrics, APE and PIE % that assess the performance of FMTS from accuracy and robustness perspectives and aid in assigning ratings. (3) A ‘data to predictions’ and ‘predictions to ratings’ workflow for rating FMTS using both uni-modal (numerical) and multi-modal (line plots (image)) data.
Perturbations: We introduce one syntactic (STP) and two semantic perturbations (SMP) (Figure 3(a)) inspired by real-world applications in unintended scenarios to assess the accuracy and robustness of FMTS. SMPs alter data meaning while preserving context, e.g., a stock’s value might change due to data entry errors or market-specific catalysts. STPs modify the structure of the data without altering the content. Drop-to-zero (P1) is an SMP inspired by common data entry errors (Ley et al. 2019). Every $n^{th}$ value in the original stock price data is set to zero. Sampling the time series with a sliding window of size $n$ ensures each sample contains a zero. In Value halved perturbation (P2), an SMP, we reduce every $n^{th}$ number in the original stock price data to half of its value. This perturbation simulates periodic adjustments, possibly reflecting events like stock splits or dividend payments. Missing values perturbation (P3), an STP, converts every $n^{th}$ number in the original stock price data to a null value, simulating real-world missing data points in financial datasets due to system incomplete transmissions.
4.1 Evaluation Metrics
@@
In this section, we describe our evaluation metrics for measuring forecasting accuracy and robustness. [^1] Full formulas are provided in the Appendix C.
Forecasting Accuracy Metrics We evaluate forecasting accuracy using three metrics (Makridakis, Spiliotis, and Assimakopoulos 2022): Symmetric mean absolute percentage error (SMAPE) measures average percentage error between actual and predicted values. Mean absolute scaled error (MASE) compares the mean absolute error of forecasts to a naive one-step forecast. Sign Accuracy quantifies how well predicted forecasts align with recent observed values. For SMAPE and MASE, lower values indicate better performance, while higher values are better for Sign Accuracy.
[…]
@@
Average Perturbation Effect (APE) Average Treatment Effect provides the average difference in outcomes between between treated and untreated units (Wang, Nianogo, and Arah 2017). In our context, it computes the difference between perturbed data residuals (P1 through P3) and the unperturbed data residuals (P0), thereby measuring the impact of the perturbation on the outcome. Hence, we refer to this metric as APE. This metric helps us answer RQ3. It is formally defined using the following equation:
$[|E[R^{max}_{t}=j|do(P=i)]-E[R^{max}_{t}=j|do(P=0)]|]$ (1)
@@
Propensity Score Matching - Deconfounding Impact Estimation % (PSM-DIE % or PIE %) In (Lakkaraju, Srivastava, and Valtorta 2024), a linear regression model was used to estimate causal effects, assuming a linear relationship between variables. This method, however, doesn’t capture non-linear relationships or fully eliminate confounding biases. They proposed the DIE % metric for binary treatment values. Our work uses six treatment (perturbation) values, applying Propensity Score Matching (PSM) (Rosenbaum and Rubin 1983) to target confounding effects by matching treatment and control units based on treatment probability, similar to RCTs and independent of outcome variables (Baser 2007). We modify DIE % to introduce PIE % (PSM-DIE%) to answer RQ2. It is defined as:
$\left[||APE_{o}|-|APE_{m}||\right]*100$ (2)
@@
4.2 Workflow
Our proposed workflow consists of two parts: Data to Predictions and Predictions to Ratings. In the first part, as shown in Figure 3(a) [^2] All the figures in the paper can be found in a higher resolution in the Appendix D., FMTS process the input and predict the next ‘d’ timesteps. The FMTS and baseline models are described in Section 5.1. In the second part, illustrated in Figure 3(b), we adapt the method from (Lakkaraju, Srivastava, and Valtorta 2024) to rate text-based sentiment analysis systems for bias. We extended their approach to handle our more complex multi-modal data with multiple perturbations, beyond the original textual data and binary treatments. The modified metrics, APE and PIE%, defined in Section 4.1, help manage this complexity. The rating algorithms are detailed in Appendix A. These metrics, referred to as raw scores, establish a partial order for determining final system ratings, which vary by rating level, L.
@@ 4.2 Workflow
Figure 3: (a) Black arrows denote the unperturbed and red arrows indicate the perturbed paths. Dashed lines shows the multi-modal path. The perturbed parts of the plots are highlighted in red. (b) Workflow for performing statistical and causal analysis to compute raw scores and assign final ratings to the test systems
5 Experiments and Results
This section introduces the FMTS used in our experiments, baseline models, test data, and evaluation metrics, including two new metrics for perturbations and confounders. We also present the user study design, responses, and findings.
5.1 Experimental Apparatus
FMTS
We used four FMs in a zero-shot setting: TS forecasting FMs (MOMENT and Chronos) and general-purpose (GP) multimodal FMs (Gemini-V and Phi-3) adapted for TS forecasting. We set $n=80$ and $d=20$. Table 1 provides an overview of the FMTS architectural details.
Model Mode Size Purpose & Arch. Inf. Time (sec/sample)
Gemini 1.5 Flash Multi 32B GP-1A, 1B, Decoder 1.6 (1A); 10.2 (1B)
Phi-3-vision Multi 4.2B GP-2A, 2B, Enc-Dec 19.7 (1A); 26.6 (1B)
MOMENT-large Uni 385M TS-1, Encoder 0.315
Chronos-T5-small Uni 46M TS-2, Enc-Dec 0.811
Table 1: Overview of the architectural details of FMTS. *Best guess in the absence of official information.
1. MOMENT ($S_{m}$) (Goswami et al. 2024) is an open-source FM that handles a variety of tasks such as forecasting, classification, anomaly detection, and imputation in zero-shot and few-shot settings, and it can also be fine-tuned if needed. We use MOMENT-large for our experiments, which utilizes T5-Large encoder (Raffel et al. 2020) as the base architecture.
2. Chronos ($S_{c}$) (Ansari et al. 2024) is a pretrained probabilistic time series model. It tokenizes time series values using scaling and quantization, then trains transformer-based language models via cross-entropy loss. Chronos uses the T5 encoder-decoder architecture, specifically Chronos-T5-Small for our experiments.
3. Gemini-V (Gemini 1.5 Flash) ($S_{g}$, $S_{g}^{ni}$(Team et al. 2023) is an FM under the Gemini series. This model is designed to understand inputs that include text and images, enabling it to generate relevant text responses based on a comprehensive analysis of the combined data. This model processes text and images to generate relevant text responses from combined data. We call the ‘numeric_only’ mode, $S_{g}$, and the ‘numeric_and_vision’ mode, ($S_{g}^{ni}$).
4. Phi-3 (Phi-3-vision-128k-instruct) ($S_{p}$, $S_{p}^{ni}$) (Abdin et al. 2024) is a lightweight, SOTA open multimodal model. It is built upon datasets that include synthetic data and filtered publicly available websites, with a focus on very high-quality, reasoning-dense data in both text and vision. We call the ‘numeric_only’ mode, $S_{p}$, and the ‘numeric_and_vision’ mode, ($S_{p}^{ni}$). Below is the prompt template we used for time-series forecasting using Gemini-V and Phi-3 models (we omit the text highlighted in blue for uni-modal forecasting:
You are a time series forecasting model that only outputs the forecasted numerical values. Input: <time series> Given the input time series for the past 80 time steps and the corresponding time series plot, can you forecast the next 20 time steps? Provide a list of 20 numeric values only. Do not provide any discussion.
Baselines
We consider the following baselines:
1. AutoRegressive Integrated Moving Average (ARIMA) ($S_{a}$) is a widely used statistical approach for time series forecasting. It combines three different components: Autoregressive (AR), differencing (I), and moving average (MA) to capture the patterns in the time-series data and predict the next ‘d’ (from section 3) values.
2. Biased system ($S_{b}$) is an extreme baseline biased towards META and GOOG (technology companies), assigning residuals of 0 and 200 respectively, while assigning higher residuals to other companies, representing maximum bias.
3. Random system ($S_{r}$) assigns random price predictions within a company range for contextually meaningful values.
Test Data
5.2 Experimental Evaluation
Figure 4: Studying each metric with respect to impact of company and industry as confounders for all models and all perturbations. Plotted in double logarithmic scale, lower values indicate better robustness. Ratings generated by our method (with $L=3$) are shown on the top of each plot. The complete final order (with ratings) are shown in Table 2 in Appendix E.
In this section, we describe the experimental setup used to address the RQs stated in Section 3, the results obtained, and the conclusions drawn from the results. Figure 2 shows the causal diagrams used to answer the RQs.
RQ1: Does Sensitive Attribute affect the Residual, even though Sensitive Attribute has no effect on Perturbation? Setup: In this experiment, the causal link from the Sensitive Attribute to Perturbation is absent, as the perturbation to the stock prices does not depend on the corresponding company name or the industry i.e., perturbations are applied uniformly across all the data points. We quantify the statistical bias exhibited by the systems by using WRS described in Section 4.1. We perform two different analyses in this experiment: one to measure the discrepancy shown across different industries (WRSIndustry) and another to measure the discrepancy among both the companies (WRSCompany) within the same industry. Results and conclusion: From Figure 4, most discrepancies can be observed across industries compared to the discrepancies across companies within each industry. When the input data was subjected to perturbation P3 ($WRS_{avg}$ of 5.35333Table 4 in Appendix E shows the RQs along with the average values for all metrics across different systems and perturbations.), the systems exhibited more statistical bias. From Figure LABEL:fig:radar-agg, $S_{a}$ ($WRS_{avg}$ of 3.96) exhibited the least statistical bias, while $S_{p}$ ($WRS_{avg}$ of 6.28) exhibited the highest statistical bias among the systems evaluated under the perturbations considered. Hence, we conclude that Sensitive Attribute affects the Residual, even though Sensitive Attribute has no effect on Perturbation.
Setup: In this experiment, we use PIE % defined in equation 2 to compare the APE (defined in equation 1) before and after deconfounding using the PSM technique as the presence of the confounder opens a backdoor path from Perturbation to Residual through the Confounder. The causal link from Confounder to Perturbation will be valid only if the perturbation applied depends on the value of the confounder (i.e. the company or the industry the specific data points belong to). To ensure the probability of perturbation assignment varies with respect to the Confounder across three distributions (DI1 through DI3) in the case of Industry and six different distributions in the case of Company (DC1 through DC6), we implement weighted sampling. For each distribution, weights are configured so that perturbation groups P1 through P3 have a twofold higher likelihood of selection compared to P0 for specific values of the confounder. For example, META in DC1, GOOG in DC2, and so on. This strategy highlights significant cases, although other combinations are possible for further exploration. Results and Conclusion: Figure 4 shows that selecting Company as the confounder leads to greater confounding bias in the systems. In Figure LABEL:fig:radar-agg, $S_{g}^{ni}$ ($PIE_{avg}\%$ of 523.09) exhibited the least confounding bias, while $S_{p}$ ($PIE_{avg}\%$ of 2677.62) exhibited the most. Systems showed more confounding bias under perturbation P3 ($PIE_{avg}\%$ of 1563.94). Therefore, the Confounder affects the relationship between Perturbation and Residual, particularly when the Confounder influences the Perturbation.
RQ3: Does Perturbation affect the Residual when Sensitive Attribute may have an effect on Residual? Setup: The experimental setup in this experiment is same as that for answering RQ2. To compute the APE, we used PSM described in Section 4.1. PSM allows us to effectively determine the effect of Perturbation on the Residual. For instance, if two matched points belong to the same company but only one was perturbed, any difference in their residuals can be directly attributed to the perturbation itself rather than to other confounding factors. This method provides a clear understanding of the true impact of the Perturbation on the Residual. As our rating method aims to bring the worst possible behavior of the systems, we take the MAX(APE) as the raw score that is used to compute the final ratings. Results and Conclusion: It is undesirable to have a higher APE, as it implies that the perturbation applied can have a significant impact on the residuals of different systems. From Figure 4, when Industry was considered as the confounder, it led to a higher APE. As the outcome of $S_{b}$ depended on the Company (and varied from one company to another), the perturbation did not have any effect on the system. Whereas, when Industry was considered as the confounder, the perturbation appeared influential, resulting in a high APE for $S_{b}$. From Figure LABEL:fig:radar-agg, perturbations had the least impact on $S_{g}^{ni}$ ($APE_{avg}$ of 6.49) and highest impact on $S_{p}$ ($APE_{avg}$ of 28.53). Among all the perturbations, P1 ($APE_{avg}$ of 21.11) was the most disruptive. Hence, Perturbation affects the Residual when Sensitive Attribute may have an effect on Residual.
RQ4: Does Perturbations degrade the accuracy of S? Setup: For this experiment, we do not use any causal model. We compute the three accuracy metrics widely used in for the task of financial time-series forecasting (Makridakis, Spiliotis, and Assimakopoulos 2022) which were summarized in Section 4.1. Results and Conclusion: From Figure LABEL:fig:radar-agg, $S_{c}$ exhibited the highest amount of forecasting accuracy in terms of SMAPE (average of 0.05) and MASE (average of 4.67), while $S_{a}$ outperformed all other systems in terms of sign accuracy (average of 58.57). $S_{b}$ consistently predicted the correct directional movement of stock prices, exhibiting high sign accuracy as it was designed to adjust residuals based on specific company stock prices. Perturbations P2, P1, and P0 caused the highest decline in SMAPE (average of 0.08), MASE (average of 8.99), and sign accuracy (average of 49.97), respectively. Hence, Perturbations degrade the accuracy of S.
5.3 Overall Performance Comparison
Now, we provide an overall comparison of the different systems across all metrics to highlight key findings about their performance under various perturbations. Figure LABEL:fig:radar-agg shows radar plots with mean forecasting accuracy and robustness metrics, including values under perturbation P2.
1. Clear Domination Signals From Figure LABEL:fig:radar-agg and the detailed results from Section 5.2, we can draw the following conclusions: $S_{c}$’s Superiority in Forecasting Metrics: $S_{c}$ consistently outperformed other models in terms of forecasting accuracy metrics, specifically SMAPE and MASE. This indicates that $S_{c}$ is highly effective in predicting stock prices with minimal error. General Superiority Over Biased and Random Systems: All models perform better than the biased and random systems in forecasting metrics. This underscores the importance of using well-designed models over naive or biased approaches. Robustness in PIE % and APE Metrics: According to the average scores, the $S_{g}^{ni}$ system demonstrated superior robustness in PIE % and APE metrics. This suggests that $S_{g}^{ni}$ is more resilient to perturbations and confounding biases compared to other systems.
2. Role of Modality The results from Figure 5 indicate that multimodal FMTS, $S_{g}^{ni}$ and $S_{p}^{ni}$, generally perform better in terms of both robustness and accuracy, suggesting that incorporating multiple data modalities (e.g., numerical and image) can improve the system’s ability to make accurate predictions and remain robust against various perturbations.
Figure 5: Effect of the modalities for $S_{g}$ (left) and $S_{p}$ (right).
3. Role of Confounders Our analysis (Figs. 4 and LABEL:fig:radar-agg: left) shows that using industry as a confounder introduces more bias, with higher PIEscores indicating significant industry-specific effects on the relationship between perturbations and residuals. Inter-industry comparisons also show more discrepancies, as evidenced by WRS scores. Conversely, the impact of company as a confounder varies. For instance, system $S_{b}$ shows minimal effect from company-specific perturbations, suggesting it is well-tailored to company data, while other systems have higher APE scores, indicating a significant impact on residuals due to company-specific factors.
4. Role of Architecture Our evaluation (Fig. 6, left) indicates that the Time Series architecture generally performs better across several metrics, such as achieving the best values in APE_C, PIE_C, SMAPE, MASE, and WRS_I, suggesting that the TS architecture may be more effective for these specific tasks compared to the general purpose architectures. Fig. 6 (right) shows that decoder-only architecture outperforms others in terms of both accuracy and robustness.
Figure 6: Role of architecture in forecasting accuracy and robustness. Performance is averaged across models within each category. See Table 1.
Overall, our comparison highlights $S_{c}$’s forecasting accuracy, the robustness of multimodal systems against perturbations and confounding biases, and the superiority of well-designed models over naive approaches.
5.4 User Study
We conducted a user study to evaluate the ratings generated by our approach for comparing the behavior of various FMTS based on two key metrics: robustness and statistical fairness (defined as lack of statistical bias). To simplify the evaluation for participants, we converted the generated ratings into rankings (i.e., the system with the highest robustness ranking is the most robust system). The main objective of this study was to validate the following hypotheses:
This IRB-approved study [^4] Institutional details anonymized for reviewing. involved participants evaluating FMTS models forecasting stock prices for companies in various industries, participants were introduced to key concepts including robustness, fairness, and error metrics (maximum residual, mean, and standard deviation of errors) to ensure informed evaluations. All inputs were sought on Likert 5-value scale. The details and complete set of questions are provided in the Appendixes  FLABEL:sec:appendix-study-q.
The study was structured into four panels: a self-assessment on knowledge about time series and financial data, a fairness panel, where fairness was the evaluated metric, and two robustness panels, where robustness was assessed under two different perturbations (P1 and P2). In the fairness panel, participants were presented with graphs depicting the residual values of six different systems and an ideal system using stock price data from the Technology and Pharmaceuticals industries. They were provided with the mean and standard deviation of errors and asked to rank the systems from least to most fair. Participants then rated the difficulty of this ranking task (1 being the most difficult). Subsequently, they were shown the rankings generated by our approach and asked to rate the accuracy of these rankings (1 being the least accurate). Finally, participants were asked to rank the difficulty of comparing the behavior of different systems using our rankings (1 being the most difficult). For the robustness panels, similar questions were posed, with users evaluating systems based on their robustness to different perturbations (P1 and P2). A total of 26 users from academia and industry participated over two weeks. We performed different types of statistical tests to draw significant conclusions; see details in Table 5, 6 in Appendix F. We now discuss the key findings from the tests.
To evaluate HP1, we conducted a paired t-test to compare user responses on difficulty of ranking various systems before and after presenting our rankings. The same participants assessed the difficulty using both the graph representation of fairness and our ranking representation, making the paired t-test appropriate. Paired t-test also accounts for the inherent correlation between the paired rankings, making it suitable to account for potential different perceptions across the two representations. Paired t-tests for each robustness panel indicated a significant difference before (P1:$\mu$ = 2.70, $\sigma$ = 1.06; P2: $\mu$ = 2.65, $\sigma$ = 1.17) and after (P1: $\mu$ = 3.23, $\sigma$ = 1.42; P2: $\mu$ = 3.07, $\sigma$ = 1.44) our rankings were presented with P1:t(26) = -1.89, p = 0.030, and P2: t(26) = -1.62, p = 0.059. Since the p-values $\prec$ 0.1, we confirm HP1 that the ranking generated by our approach significantly reduced the perceived difficulty of comparing different systems. Same approach is used to evaluate HP2. The paired t-test showed no significant change in perceived difficulty scores before ($\mu$ = 2.54, $\sigma$ = 1.30) and after ($\mu$ = 2.92, $\sigma$ = 1.35) our rankings were presented, t(26) = -1.18, p = 0.12. Since p $\succ$ 0.1, we conclude that our ranking representation did not significantly reduce the perceived difficulty of comparing different systems. The lack of significant reduction in perceived difficult may have stemmed from the complexity of the graphical representations.
To validate HP3, we used the Spearman Rank Correlation coefficient (Zar 2005) ($\rho$) to evaluate the alignment between the users’ rankings and those produced by our approach. We considered a confidence interval of 90 %. The fairness panel showed a high correlation ($\rho$= 0.73), and the robustness under P1 showed a strong correlation ($\rho$ = 0.91). However, robustness under P2 showed a weak correlation ($\rho$ = 0.14).
In summary, the results of the user study indicate that the rankings generated by our approach can significantly reduce the difficulty of comparing the robustness of different FMTS systems. However, when the comparison metric is fairness, this reduction is not significant. Additionally, while user rankings align well with our method generated rankings for fairness and robustness under P1, they show a weak correlation for robustness under P2, indicating differing perceptions of the ’value halved’ perturbation. P1 (drop-to-zero) involves a significant semantic change that is easier to spot, whereas P2 (value halved) is also a semantic perturbation but subtler, making it potentially harder to identify. Our current study is preliminary and promising; an avenue for future work is to conduct it at a larger scale.
6 Discussion and Conclusion
Our paper aimed to measure the impact of perturbations and confounders on the outcome of FMTS using stock prices across leading companies and industries. We studied Industry and Company as confounders, motivated by the intuition that stakeholders rely on learning-based systems for stock purchase decisions and would be interested in knowing if model errors depend on stock price ranges. For example, does a model commit more errors predicting META’s stock prices compared to MRK’s? To minimize volatility effects, we performed both intra-industry and inter-industry analyses. In future, we plan to study confounders like seasonal trends and financial news. As demonstrated, we believe metrics should be selected based on the questions one wants to answer, rather than relying solely on statistical accuracy. The hypothesis testing approach from (Lakkaraju, Srivastava, and Valtorta 2024), adapted for our work, helped quantify biases and perturbation impacts on test systems. The perturbations used in our analysis have real-world impacts, applicable to both numeric and multi-modal data. While methods like differential evaluation can find the most impactful perturbation variations, we focused on assessing whether simple, subtle perturbations affect FMTS.
Conclusion We proposed a causally grounded empirical framework to study FMTS robustness against three input perturbations, evaluating seven state-of-the-art FMTS across six prominent stocks in three industries. Our framework’s ratings accurately assessed FMTS robustness and provided actionable insights for model selection and deployment. Experiments showed multi-modal FMTS exhibited greater robustness, while uni-modal FMTS had higher forecasting accuracy. FMTS trained on time series tasks showed better robustness and accuracy compared to general-purpose FMTS. A user study confirmed our ratings effectively convey FMTS robustness to end-users, demonstrating the framework’s real-world applicability.
Appendix A Details of Rating Algorithms
Apart from Algorithm 2, the rest were adapted from (Lakkaraju, Srivastava, and Valtorta 2024) to suit the FMTS forecasting setting. Here is how the rating method works.
  1. 1.
    Algorithm 1 computes the weighted rejection score (WRS) which was defined in Appendix C in the main paper.
  2. 2.
    Algorithm 2 computes the PIE % based on Propensity Score Matching (PSM) which was defined in Section 4.1 in the main paper.
  3. 3.
    Algorithm 3 creates a partial order of systems within each perturbation based on the raw scores computed. It will arrange the systems in ascending order w.r.t the raw score. The final partial order (PO) will be a dictionary of dictionaries.
  4. 4.
    Algorithm 4 computes the final ratings for systems within each perturbation based on the PO from previous algorithm. It splits the set of raw score values obtained within each perturbation into ‘L’ parts where ‘L’ is the rating level chosen by the user. Each of the systems is given a rating based on the compartment number in which its raw score belongs. The algorithm will return a dictionary with perturbations as keys and ratings provided to each system within the perturbation as the value.
@@ Appendix A Details of Rating Algorithms
$d$, dataset corresponding to a specific perturbation.
$CI$, confidence intervals (95%, 70%, 60%).
$s$, system for which WRS is being computed.
$W$, weights corresponding to different CIs (1, 0.8, 0.6).
@@
$Z$, Sensitive attribute
$\psi$, weighted rejection score. $\psi\leftarrow 0$ for each $ci_{i},w_{i}\in CI,W$ do
       // $z_{m},z_{n}$ are classes of $Z$
      for each $z_{m},z_{n}\in Z$ do
             $t,pval,dof\leftarrow T-Test(z_{m},z_{n})$;
@@
Purpose: is used to calculate the Deconfounding Impact Estimation using Propensity Score Matching (PIE).
@@
$s$, a system belonging to the set of test systems, $S$.
$D$, datasets pertaining to a perturbation (different distributions).
$p$, A perturbation other than $p_{0}$
$p_{0}$, control perturbation (or no perturbation.
@@
$\psi\leftarrow 0$
$PIE\_list\leftarrow[]$             // To store the list of PIE % of all the datasets.
for each $d_{j}\in D$ do
       $APE\_o\leftarrow E(R|P=p_{m})-E(R|P=p_{0})$;
       $APE\_m\leftarrow E(R|do(P=p_{m}))-E(R|do(P=p_{0}))$;
       $PIE\_list[j]\leftarrow(APE\_m-APE\_o)*100$;
end for
$\psi\leftarrow MAX(PIE\_list)$;

### v3

@@ 4. Solution Approach
Our solution approach comprises the following components: (1) A set of six perturbations applied selectively to numerical time series data, time-frequency spectrograms, time-series intensities, and time-series line plots. (2) Robustness evaluation metrics: WRS, APE, and PIE %, and forecasting accuracy metrics, each used separately to compute the ratings. (3) A structured workflow that maps data to predictions and predictions to ratings, enabling the evaluation of TSFM using both uni-modal (numerical) and multi-modal (line plots, time-frequency, and time-intensity data combined with numerical data) inputs.
4.1. Perturbations
Figure 2. Unperturbed image and its perturbed variations, along with a corresponding difference image that highlights (in red) the portion that is modified after perturbation. In (f,g), the single pixel change is highlighted with a white circle around it.
We define six perturbations: two semantic (SP), one syntactic (SyP), two input-specific (IP), and one composite (CP) (Figure 2) inspired by real-world applications in unintended scenarios to assess the robustness and accuracy of TSFM.
a) Semantic Perturbation (SP): Semantic perturbations are alterations made to data that change its meaning while preserving the overall context. For example, in time-series forecasting, a stock’s value might change drastically due to errors in data entry, or it might fluctuate due to some market-specific catalyst that affects certain companies. Under SP, we consider two perturbations:
1. Drop-to-zero (P1): It is inspired by common data entry errors (Ley et al., 2019). Every $n^{th}$ value in the original stock price data is set to zero. Sampling the time series with a sliding window of size $n$ ensures each sample contains a zero.
2. Value halved perturbation (P2): Every $n^{th}$ value in the original stock price data is reduced to half of its value. This perturbation simulates periodic adjustments, possibly reflecting events like stock splits or dividend payments.
b) Syntactic Perturbation (SyP): SyPs modify the structure of the data without altering its fundamental meaning. We consider one such perturbation.
1. Missing values perturbation (P3) Every $n^{th}$ value in the original stock price data is converted to a null value, simulating real-world missing data points in financial datasets due to incomplete transmissions.
c) Input-specific Perturbation (IP): Input-specific perturbations are alterations specific to the mode - features and context of the data being used. In time-series forecasting, altering some pixels (e.g., changing their color) in a time-frequency spectrogram image is one example. Under IP, we consider two different perturbations:
1. Single pixel change (P4): In (Su et al., 2019), the authors modified a single pixel in each of the test images. With this approach, they fooled three types of DNNs trained on the CIFAR dataset. In our work, we alter the center pixel of each multi-modal input to black based on the intuition that small and consistent change to the images can significantly alter the models’ predictions.
2. Saturation change (P5): In (Zhu et al., 2023), the authors showed that the adversarial perturbations in the S-channel (or saturation channel) of an image in HSV (Hue Saturation Value) form ensures a high success rate for attacks compared to other channels. In our work, we increase the saturation of the multi-modal input ten-fold based on the intuition that a subtle change can affect the models’ predictions.
d) Composite Perturbation (CP or P6): We consider a composite case where TSFM is combined with another AI system to reflect the influence of market sentiment on stock prediction (Mishev et al., 2020). We assess sentiment for each time series by passing the corresponding time-series line plot to a zero-shot CLIP-based sentiment analysis system (SAS) (Radford et al., 2021; Bondielli et al., 2021), which outputs sentiment intensity values (negative (-1), neutral (0), positive (1)). These labels are scaled to [0, 255] and represented as a sentiment intensity stripe. This zero-shot CLIP-based SAS may exhibit bias or inaccuracies; our goal is to study the robustness of MM-TSFM in conjunction with such an AI system.
4.2. Evaluation Metrics
@@
In this section, we describe our evaluation metrics for measuring robustness and forecasting accuracy.
4.2.1. Robustness Metrics
We adapt the Weighted Rejection Score (WRS) originally proposed in (Lakkaraju et al., 2024) to measure statistical bias. Additionally, we introduce two new metrics: APE and PIE % (modified versions of ATE (Abdia et al., 2017) and DIE % (Lakkaraju et al., 2024)) tailored to answering our research questions:
Weighted Rejection Score (WRS): WRS quantifies statistical bias across protected attributes ($Z$) by assessing the extent to which the null hypothesis is rejected at various confidence intervals (CIs). For a given protected attribute $z_{i}\in Z$, outcome distributions $(R^{max}_{t}|z_{i})$ are compared across all pairs of groups within $z_{i}$ (for e.g., if $z_{i}$ has $m$ protected groups, $^{m}C_{2}$ pairwise comparisons are performed). The t-value for each pair is computed using Student’s t-test (Student, 1908). Let $x_{i}$ represent the number of rejections for different CI, and $w_{i}$ be the weight assigned to that CI. Specifically, weights of 1, 0.8, and 0.6 are used for 95 %, 75 %, and 60 % CIs, respectively. WRS is given by:
(1) $WRS=\sum_{i}w_{i}*x_{i}$
@@ 4.2.1. Robustness Metrics
Average Perturbation Effect (APE): In causal inference, Average Treatment Effect (ATE) provides the average difference in outcomes between between treated and untreated units (Wang et al., 2017). In our context, it computes the difference between perturbed data residuals (P1 through P6) and the unperturbed data residuals (P0), thereby measuring the impact of the perturbation on the outcome. Hence, we refer to this metric as APE. It is formally defined using the following equation:
(2) $|E[R^{max}_{t}=j|do(P=i)]-E[R^{max}_{t}=j|do(P=0)]|]$
@@
Propensity Score Matching - Deconfounding Impact Estimation % (PSM-DIE % or PIE %) In (Lakkaraju et al., 2024), a linear regression model was used to estimate causal effects, assuming a linear relationship between variables. This method, however, doesn’t capture non-linear relationships or fully eliminate confounding biases and only works for binary treatments. Our work uses six treatment (perturbation) values, applying Propensity Score Matching (PSM) (Rosenbaum and Rubin, 1983) to target confounding effects by matching treatment and control units based on treatment probability, similar to RCTs and independent of outcome variables (Baser, 2007). It is defined as:
(3) $\left[||APE_{o}|-|APE_{m}||\right]*100$
@@
4.2.2. Forecasting Accuracy Metrics
We evaluate forecasting accuracy using three metrics (Makridakis et al., 2022):
Symmetric mean absolute percentage error (SMAPE) measures the relative difference between predicted and actual values and assigns equal weight to over- and under-estimations. It is defined as,
(4) $SMAPE=\frac{1}{T}\sum_{t=1}^{T}\frac{|x_{t}-\hat{x}_{t}|}{(|x_{t}|+|\hat{x}_{t}|)/2},$
(5) $MASE=\frac{\frac{1}{T}\sum_{i=t+1}^{t+T}|x_{i}-\hat{x_{i}}|}{\frac{1}{t}\sum_{i=1}^{t}|x_{i}-x_{i-1}|},$
Sign Accuracy quantifies the proportion of correctly predicted directional changes in the time series. A higher accuracy indicates better alignment with actual trend movements.
(6) $\text{Sign Accuracy}=\frac{1}{T}\sum_{t=1}^{T}1\left(\text{sign}(\hat{x}_{t}-\hat{x}_{t-1})=\text{sign}(x_{t}-x_{t-1})\right),$
where $T$ is the total number of predicted time steps, $x_{t}$ is the actual value, and $\hat{x}_{t}$ is the predicted value at time $t$. The indicator function $1(\cdot)$ returns 1 if the predicted sign matches the actual sign and 0 otherwise. Higher Sign Accuracy values indicate better directional prediction.
4.3. Workflow
Our proposed workflow consists of two components: Data to Predictions and Predictions to Ratings. In the first stage, as shown in Figure 3(a), TSFM processes the input and predicts the next ‘d’ timesteps. The TSFM and baseline models are detailed in Section 5.1.1. In the second stage, illustrated in Figure 3(b), we extend the approach from (Lakkaraju et al., 2024), originally designed for assessing SASs, to accommodate our more complex multi-modal data with multiple perturbations, beyond the original textual data and binary treatments. The modified metrics, APE and PIE %, introduced in Section 4.2, help manage this complexity. These raw scores establish a partial order for determining final model ratings, which vary based on the rating level, $L$. The following four algorithms provide the detailed implementation of metrics, raw score calculation, and the final ratings calculation:
  • Algorithm 1 computes WRS, as defined in Section 4.2, to measure statistical bias by analyzing how sensitive the model’s outcomes are to variations in sensitive attributes. This helps us answer RQ1 from Section 3.
  • Algorithm 2 calculates the PIE %, as defined in Section 4.2, to assess confounding bias by measuring the effect of the confounder on model outcomes before and after deconfounding. This can help us answer RQ2 from Section 3.
  • Algorithm 3 calculates APE by evaluating the difference in the model’s outcomes for perturbed data and unperturbed data to evaluate the impact of the perturbation on the outcome as defined in Section 4.2. This helps us answer RQ3 from Section 3.
  • Algorithm 4 generates a partial order of TSFM for each perturbation based on their WRS or PIE or ATE scores. Models are ranked based on these raw scores, and the rankings are stored in a dictionary mapping each perturbation to its corresponding system order (as shown in Tables 5 and 6).
  • Algorithm 5 assigns final ratings to systems for each treatment based on the partial order generated by Algorithm 3. It partitions the raw scores within each treatment into $L$ user-defined rating levels and assigns ratings accordingly. The output is a dictionary where treatments serve as keys, mapping to model ratings as values. Lower ratings indicate better performance and greater robustness, except for Sign Accuracy, where higher ratings correspond to better performance.
@@ 4.3. Workflow
Figure 3. (a) Workflow illustrating the data modalities, the perturbations applied to each, and their propagation through different models to generate predictions. P0 is unperturbed, P1 (Drop-to-zero) and P2 (Value halved) are semantic perturbations (SMP), P3 (Missing Values) is a syntactic perturbation (SNP), P4 (Single pixel change) and P5 (Saturation change) are input-specific perturbations (IP), and P6 is a composite perturbation (CP). (b) Workflow for performing statistical and causal analysis to compute raw scores and assign final ratings to the test systems.
@@
$d$, dataset corresponding to a specific perturbation; CICI, confidence intervals (95%, 70%, 60%); $s$, a model belonging to the set of test models, $S$; $W$, weights corresponding to different CIs (1, 0.8, 0.6).
@@
$Z$, Sensitive attribute; $\psi$, weighted rejection score.
$\psi\leftarrow 0$
for each $ci_{i},w_{i}\in CI,W$ do
  // $z_{a},z_{b}$ are classes of $Z$
 for each $z_{a},z_{b}\in Z$ do
     $t,pval,dof\leftarrow T-Test(z_{a},z_{b})$;
@@
Purpose: Calculate the Deconfounding Impact Estimation (PIE) using Propensity Score Matching (PSM).
@@
$s$, $d$ (as defined in the previous algorithm); $p$, a perturbation other than $p_{0}$ (control or no perturbation).
@@
$APE\_o\leftarrow E(R^{max}_{t}|P=p)-E(R^{max}_{t}|P=p_{0})$      // Observational APE.
$APE\_m\leftarrow E(R^{max}_{t}|do(P=p))-E(R^{max}_{t}|do(P=p_{0}))$      // Deconfounded APE.
$\psi\leftarrow(APE\_m-APE\_o)*100$      // Compute PIE.